# Retrieval-Augmented Decision Transformer: External Memory for In-context RL

### v2

@@ B.5 Procgen
The Procgen benchmark consists of 16 procedurally-generated video games and was designed to test the generalization abilities of RL agents [Cobbe et al., 2020]. Unlike other environments considered in this work, Procgen environments emits $3\times 64\times 64$ images as observations. All 16 environments share a common action space of 15 discrete actions. The procedural generation in Procgen is controlled by setting an environment seed. The environments seed randomizes the background and colour of the environment, but retains the same game dynamics. This results in visually diverse observations for the same underlying task, as illustrated in Figure 12 for three seeds on the game starpilot. The rewards in Procgen can be dense or sparse depending on the environment.
We follow Raparthy et al. [2023] and use 12 tasks for training and 4 tasks for evaluation, which we refer to as PG12 and PG4, respectively. The PG12 tasks are:
@@
Figure 12: Illustration of procedural generation in Procgen starpilot. For different seeds, the same environment looks visually considerably different. We train on multi-task dataset of 12 Procgen tasks, with each dataset containing trajectories from 200 environment seeds. To test for ICL, we evaluate on single hold-out seeds.
Training Dataset. We generate datasets by training task-specific PPO agents for 25M timesteps on 200 environment seeds per task in easy difficulty, as proposed in by Cobbe et al. [2020]. We train PPO using the same hyperparameter settings as Cobbe et al. [2020], using a learning rate of $5e^{-4}$, batch size 2048, number of update epochs of 3, entropy coefficient of $0.01$, GAE $\lambda=0.95$, and with reward normalization. We use 256 timesteps per rollout over 64 parallel environments, which results in 16384 environment steps per rollout in total. Furthermore, we found it useful to decrease the discount factor to 0.99.
As in previous experiments, we record the entire replay buffer and consequently, the datasets contain mixed-quality behaviour. We subsample the 25M transitions per task, by storing only the observations of the first 5 parallel environments, which results in approximately 2M transitions per task. To ensure disk-space efficiency, all trajectories are stored in separate hdf5 files in the lowest compression level files, with all image-observations encoded in unit8. Consequently, the datasets for all 16 tasks (32M transitions) take up only 70GB of disk space, and their hdf5 format enables targeted reading from disk, without loading an entire trajectory into RAM. We release two versions of our datasets: a smaller one containing 2M transitions per task as used in our experiments, and a larger one containing 20M transitions per task.
@@ C.1 General
Training & Evaluation. We compare RA-DT against DT, AD, and DPT on all environments. On grid-world environments, we train all methods for 100K steps and evaluate after every 25K steps. For Meta-World, DMControl and Procgen, we train for 200K steps and evaluate after every 50K steps. During evaluation, the agent is given 40 interaction episodes for ICL on Dark-Room and Dark Key-Door, and 30 episodes on MazeRunner, Meta-World, DMControl, and Procgen. We use the ICL curves as the primary evaluation mechanism, and report the scores at the last evaluation step (100K or 200K). Following, Agarwal et al. [2021] we report the mean and 95% confidence intervals across tasks and over 3 seeds in all experiments.
@@
Across experiments, we keep most parameters fixed, unless mentioned otherwise. We train with a batch size of 128 on all environments, except for $40$ grids, where we use a batch size of 32. We use a constant learning rate of $1e^{-4}$ and 4000 linear warm-up steps followed by a cosine decay to $1e^{-6}$ and train using the AdamW optimizer [Loshchilov & Hutter, 2018]. Furthermore, we employ gradient clipping of 0.25, weight decay of 0.01, and a dropout rate of 0.2 for all methods.
@@
Context Length. On grid-worlds, we use a context length $C$ equivalent to two 2 episodes for AD, DPT and DT. For example, on $40\times 20$ grids, this results in a sequence length of 6400 ($=1600*4$ for state/action/reward/RTG) for the DT and a sequence length of 4800 for AD. On Meta-World, DMControl and Procgen, we reduce the sequence context length to 50 steps for DT. For RA-DT, we use a shorter context length of $C=50$ transitions across environments, except for $20\times 20$ and $40\times 20$ grids, where we increase the context length to 100. We want to highlight, that the context length for RA-DT applies to both the input context and the retrieved context. The retrieved context contains the past, and future context, as described in Section 3.2.1. Consequently, the effective context length of RA-DT is $C+2*C$ and is independent of the episode length.
@@
Network Architecture. For all environments, except for Procgen, we use a GPT2-like network architecture [Radford et al., 2019] with 4 Transformer layers, 8 head and hidden dimension of 512, which results in 16M parameters. On Procgen, we use a larger model with 6 Transformer blocks, 12 heads and hidden dimension of 768. States, actions, rewards and RTGs are embedded using separate embedding layers per modality, as proposed by Chen et al. [2021]. For all modalities and environments, we use standard linear layers to embed the inputs. Procgen is again an exception, where we use the convolutional architecture proposed by Espeholt et al. [2018] and adopted in prior works [Cobbe et al., 2020; Schmidt & Schmied, 2021; Schwarzer et al., 2023]. Processing image-sequences is computationally demanding. Therefore, we first pre-train the vision-encoder using a separate DT and embed all images in the dataset using the learned vision encoder. Therefore, the data-loading is not bottlenecked by loading entire images into memory, but only their compact representations.
@@
Furthermore, we use global positional embeddings. We also experimented with the Transformer++ recipe (RoPE, SwiGLU, RMSNorm), but only observed minimal performance gains for our problem setting. To speed-up training, we use mixed-precision Micikevicius et al. [2017], model compilation as supported in PyTorch [Paszke et al., 2019], and FlashAttention [Dao, 2023].
@@
Implementation. Our implementation of the DT is based on the transformers library [Wolf et al., 2020] and stable-baselines3 [Raffin et al., 2021]. We integrated AD, DPT, and RA-DT on top of this implementation.
@@ C.2 Decision Transformer
For Dark-Room and Dark Key-Door, we sample the target return for RTG conditioning before every episode $\mathcal{N}(90,5)$, $\mathcal{N}(370,10)$, and $\mathcal{N}(500,10)$ for grid sizes $10\times 10$, $20\times 20$, and $40\times 20$, respectively. On grid-worlds, we found that sampling the target return performs better than using a fixed target return per grid size. We assume this is, because specifying a particular target return biases the DT towards particular goal locations. For MazeRunner, we use a constant target return of 3. For Meta-World, DMControl, and Procgen, we set the target return the maximum return achieved for a particular task in the training datasets. However, we also found that constant target returns per domain work decently.
@@ C.3 Algorithm Distillation
AD obtains a context trajectory and learns to predict actions of an input trajectory taken $K$ episodes later. Therefore, we tune $K$ per domain. On grid-worlds, we found $K=100$ to perform the best, similar to Lee et al. [2023]. For MazeRunner and Meta-World, we set $K=1000$, and for DMControl and Procgen, we set $K=250$.
@@ C.4 Retrieval-Augmented Decision Transformer
Embedding Model. For the embedding model $g(\cdot)$, we either use a DT pre-trained on the same environment with the same hyperparameters as listed in Section C, or a pre-trained and frozen LM. For the pre-trained LM, we use bert-base-uncased from the transformers library by default. BERT is an encoder-only LM with 110M parameters, vocabulary size $v=30522$, and embedding dimension of $d_{\text{LM}}=768$ [Devlin et al., 2019]. We apply FrozenHopfield with $\beta=10$ to state, action, reward and RTG tokens (see Equation 2). To achieve this, we one-hot encode all discrete input tokens, such as actions in Dark-Room/MazeRunner/Procgen or states in Dark-Room, and rewards/RTGs in the sequence before applying the FH. For other tokens, such as continuous states/actions as in Meta-World/DMControl, we directly apply the FH. We evaluate other alternatives for the LM in Appendix E.
Constructing queries/keys/values. Regardless of whether $g$ is domain-specific or domain-agnostic, we obtain $C$ embedded tokens after applying $g$ to the input trajectory $\tau_{in}$. Subsequently, we apply mean aggregation over the context length $C$ to obtain the $d_{r}$-dimensional query representation. We experimented with aggregating over all tokens or only tokens of a particular modality (state/action/reward/RTG), and found aggregation over states-only to be most effective (see Appendix E.4). As described in Section 3.2.1, we construct the key-value pairs in our retrieval index by embedding all sub-trajectories in the dataset $\mathcal{D}$ using our embedding model $g$, $\mathcal{K}\times\mathcal{V}=\{(g(\tau_{i,t-C:t}),\tau_{i,t-C:t+C})\mid 1\leq i\leq|\mathcal{D}|\}$. To avoid redundancy, in practice we construct $H/C$ key-value pairs for a given trajectory $\tau$ with episode length $H$ and sub-sequence length $C$, instead of constructing the key and values for every step $t\in[1,H]$. Note that the values, we store $\tau_{i,t-C:t+C}$, contain both the sub-trajectory itself ($\tau_{i,t-C:t}$) and its continuation ($\tau_{i,t:t+C}$). Similar to Borgeaud et al. [2022], we found this choice important for high performance in RA-DT, because it allows the model to observe how the trajectory may evolve if it predicts a certain action (given that the retrieved context is similar enough).
Vector Index. We use Faiss [Johnson et al., 2019; Douze et al., 2024] to instantiate our vector index $\mathcal{I}$. This allows us to search our vector index in $\mathcal{O}(\log M)$ time using Hierarchical Navigable Small World (HNSW) graphs. However, in practice we found it faster to use a Flat index on the GPU as provided by Faiss instead of using HNSW, because our retrieval datasets are small enough. We use retrieval both during training and during inference. It is, however, possible to pre-compute the retrieved trajectories for $\mathcal{D}$ prior to the training phase to limit the computational demand of retrieval, as suggested by Borgeaud et al. [2022]. During evaluation, we can retrieve after every environment step or only after every $t$ environment steps. Here, $t$ represents a trade-off between inference time and final performance. We use $t=1$ for Dark-Room and Dark Key-Door, and $t=25$ for all other environments (see Appendix E.6 for an ablation on this design choice). For all environments, except for Meta-World and DMControl, we provide a single retrieved sub-trajectory in the agent’s context. For Meta-World and DMControl, we found that providing more than one retrieved sub-trajectory benefits the agent’s performance. Therefore, for these two environments, we retrieve the top-4 sub-trajectories, order them by return achieved in that trajectory, and provide their concatenation as retrieved context for RA-DT.
Reweighting. To implement the reweighting mechanism, as described in Section 3.2.3, we first retrieve the top $l\gg k$ experiences and the select the top-$k$ experiences according to their reweighted scores. We set $l=50$ in all our experiments.
Embedding Retrieved Context. After the most similar trajectories have been retrieved, we embed the state/action/reward/RTG tokens with a separate embedding layers (as is done for the regular input sequence) before incorporating them via the CA layers. We also experimented with sharing/detaching the regular embedding layers, but found it most effective to maintain separate ones. Furthermore, we experimented with an additional Transformer-based encoder for the retrieved sequences, as proposed by Borgeaud et al. [2022], but did not observe substantial performance gains despite increased computational cost.
One simple strategy to mitigate this issue is deduplication, i.e., to discard duplicate experiences before the training phase of RA-DT. To achieve this, we first construct our index as described in Section 3.2. For every key $\mathbf{k}\in\mathbf{K}$, we retrieve the top-$k$ neighbours (excluding experiences from the same episode as $\mathbf{k}$). If the similarity score is above a cosine similarity of $0.98$, we discard the experience. This substantially reduces the number of experiences in the index and speeds-up retrieval.
(a) Dark-Room 10$\times$10
(b) Dark-Room 20$\times$20
(c) Dark-Room 40$\times$20
Figure 14: In-context learning performance on (a) Dark-Room 10$\times$10, (b) Dark-Room 20$\times$20, (c) Dark-Room 40$\times$20 at end of training (100K steps). We evaluate each agent for 40 episodes on each of the 80 training tasks and report mean reward (+ 95% CI) over 3 seeds.
Two other strategies for regularizing retrieval during the training phase, are similarity cut-off and query dropout [Yasunaga et al., 2023]. Similarity cut-off first retrieves the top $m>l$ experiences, discards the experiences with a similarity score above a threshold (e.g., $0.98$), and retains only the remaining experiences $l$. If used in combination with reweighting, we set $m=2*l$. Query dropout randomly drops-out tokens (e.g., 20%) of the embedded sub-trajectory $\tau_{\text{in}}$, which leads to more diverse retrieved experiences. We found both strategies effective for RA-DT. We use query dropout of 0.2, similarity cut-off of 0.98, and deduplication by default. Furthermore, for Meta-World and DMControl, we found query-blending useful. Query-blending interpolates between then actual query and a randomly selected key from the retrieval index, $\bm{q}^{\prime}=\bm{q}*\alpha_{\text{blend}}+(1-\alpha)\bm{q}_{\text{rand}}$. For Meta-World and DMControl we additionally set $\alpha_{\text{blend}}=0.5$.
On Dark-Room and Dark Key-Door environments, we found it useful to replace retrieved experiences with experiences randomly sampled from the same task, if the query sub-sequence is from the beginning of the episode (i.e., smaller than timestep 10). This is because on these two environments, retrieving appropriate experience can be difficult if the given query sub-sequence is too short.
@@
0:  DT $\pi_{\theta}$, embed model $g$, dataset $\mathcal{D}$, gradient steps $N$, context len $C$, batch size $B$, eval frequency $E$, loss function $\mathcal{L}$ (cross-entropy or MSE), evaluate, batch-wise procedures retrieve, reweight, and update
1:  $\mathcal{I}\leftarrow\emptyset$triangleright Initialize retrieval index $\mathcal{I}$
@@
3:     $\mathcal{I}\leftarrow\mathcal{I}\cup\{(g(\tau_{t-C:t}),\tau_{t-C:t+C})\mid t\in\text{range}(0,|\tau|,C)\}$triangleright Add k-v pairs of sub-trjs to $\mathcal{I}$
@@
6:     $\mathbf{b}\sim\mathcal{D}$ where $\mathbf{b}=\{\tau_{j}\ \mid 1\leq j\leq B\}$triangleright Sample batch of sub-trjs each of length $C$
7:     $\mathbf{q}=g(\mathbf{b})$triangleright Construct queries for all sub-trjs
8:     $\mathcal{R}\leftarrow\text{{retrieve}}(\mathbf{q},\mathcal{I})$triangleright Retrieve top-$l$ sub-trjs, Eq. 3
9:     $\mathcal{S}\leftarrow\text{{reweight}}(\mathcal{R})$triangleright Re-weight top-$k$ sub-trjs, Eq. 4, 5
10:     $\mathbf{a}=\pi_{\theta}(\cdot\mid\mathbf{b},\{\tau_{\text{ret}}\in\mathcal{S}\})$triangleright Predict actions for batch
11:     $\pi_{\theta}\leftarrow\texttt{update}(\pi_{\theta},\mathcal{L},\mathbf{a},\mathbf{b})$triangleright Perform gradient step, see Appendix C.1 for $\mathcal{L}$

### v3

@@ B.5 Procgen
The Procgen benchmark consists of 16 procedurally-generated video games and was designed to test the generalization abilities of RL agents (Cobbe et al., 2020). Unlike other environments considered in this work, Procgen environments emit $3\times 64\times 64$ images as observations. All 16 environments share a common action space of 15 discrete actions. The procedural generation in Procgen is controlled by setting an environment seed. The environment’s seed randomizes the background and colour of the environment, but retains the same game dynamics. This results in visually diverse observations for the same underlying task, as illustrated in Figure 12 for three seeds on the game starpilot. The rewards in Procgen can be dense or sparse depending on the environment.
We follow Raparthy et al. (2023) and use 12 tasks for training and 4 tasks for evaluation, which we refer to as PG12 and PG4, respectively. The PG12 tasks are:
@@
Figure 12: Illustration of procedural generation in Procgen starpilot. For different seeds, the same environment looks visually considerably different. We train on a multi-task dataset of 12 Procgen tasks, with each dataset containing trajectories from 200 environment seeds. To test for ICL, we evaluate on single hold-out seeds.
Training Dataset. We generate datasets by training task-specific PPO agents for 25M timesteps on 200 environment seeds per task in easy difficulty, as proposed by Cobbe et al. (2020). We train PPO using the same hyperparameter settings as Cobbe et al. (2020), using a learning rate of $5e^{-4}$, batch size 2048, number of update epochs of 3, entropy coefficient of $0.01$, GAE $\lambda=0.95$, and with reward normalization. We use 256 timesteps per rollout over 64 parallel environments, which results in 16384 environment steps per rollout in total. Furthermore, we found it useful to decrease the discount factor to 0.99.
As in previous experiments, we record the entire replay buffer and consequently, the datasets contain mixed-quality behaviour. We subsample the 25M transitions per task by storing only the observations of the first 5 parallel environments, which results in approximately 2M transitions per task. To ensure disk-space efficiency, all trajectories are stored in separate hdf5 files in the lowest compression level, with all image-observations encoded in unit8. Consequently, the datasets for all 16 tasks (32M transitions) take up only 70GB of disk space, and their hdf5 format enables targeted reading from disk, without loading an entire trajectory into RAM. We release two versions of our datasets: a smaller one containing 2M transitions per task as used in our experiments, and a larger one containing 20M transitions per task.
@@ C.1 General
Training & Evaluation. We compare RA-DT against DT, AD, and DPT on all environments. In grid-world environments, we train all methods for 100K steps and evaluate after every 25K steps. For Meta-World, DMControl, and Procgen, we train for 200K steps and evaluate after every 50K steps. During evaluation, the agent is given 40 interaction episodes for ICL on Dark-Room and Dark Key-Door, and 30 episodes on MazeRunner, Meta-World, DMControl, and Procgen. We use the ICL curves as the primary evaluation mechanism, and report the scores at the last evaluation step (100K or 200K). Following Agarwal et al. (2021), we report the mean and 95% stratified bootstrap confidence intervals across tasks and over three seeds in all experiments. To construct the CI bands for the learning curves, we use percentile bootstrap and 2K bootstrap samples, as suggested by Agarwal et al. (2021). In contrast to Agarwal et al. (2021), we do not report the interquartile mean in our experiments, but instead opt for the mean to report in line with prior in-context RL methods (Laskin et al., 2022; Lee et al., 2023). Note that the evaluations conducted in this work are computationally costly because of the multi-task settings and environments with long episodes. Therefore, we limit our experiments to three seeds per method. Given the large number of experiments we report in this work in over 800 tasks (3 variants of Dark-Room with 100 rooms each, 3 variants of Dark Key-Door with 100 rooms, 120 mazes for MazeRunner, 50 Meta-World tasks, 16 DMControl tasks, 16 Procgen tasks) for 5 methods, increasing the number of seeds would have exceeded our computational budget.
@@
Across experiments, we keep most parameters fixed, unless mentioned otherwise. We train with a batch size of 128 on all environments, except for $40$ grids, where we use a batch size of 32. We use a constant learning rate of $1e^{-4}$ and 4000 linear warm-up steps followed by a cosine decay to $1e^{-6}$ and train using the AdamW optimizer (Loshchilov & Hutter, 2018). Furthermore, we employ gradient clipping of 0.25, weight decay of 0.01, and a dropout rate of 0.2 for all methods.
@@
Context Length. On grid-worlds, we use a context length $C$ equivalent to two 2 episodes for AD, DPT, and DT. For example, on $40\times 20$ grids, this results in a sequence length of 6400 (=16004=1600*4 for state/action/reward/RTG) for the DT and a sequence length of 4800 for AD. On Meta-World, DMControl, and Procgen, we reduce the sequence context length to 50 steps for DT. For RA-DT, we use a shorter context length of $C=50$ transitions across environments, except for $20\times 20$ and $40\times 20$ grids, where we increase the context length to 100. We want to highlight that the context length for RA-DT applies to both the input context and the retrieved context. The retrieved context contains the past and future context, as described in Section 3.2.1. Consequently, the effective context length of RA-DT is C+2CC+2*C and is independent of the episode length.
@@
Network Architecture. For all environments, except for Procgen, we use a GPT2-like network architecture (Radford et al., 2019) with 4 Transformer layers, 8 heads, and a hidden dimension of 512, which results in 16M parameters. On Procgen, we use a larger model with 6 Transformer blocks, 12 heads, and a hidden dimension of 768. States, actions, rewards, and RTGs are embedded using separate embedding layers per modality, as proposed by Chen et al. (2021). For all modalities and environments, we use standard linear layers to embed the inputs. Procgen is again an exception, where we use the convolutional architecture proposed by Espeholt et al. (2018) and adopted in prior works (Cobbe et al., 2020; Schmidt & Schmied, 2021; Schwarzer et al., 2023). Processing image sequences is computationally demanding. Therefore, we first pre-train the vision encoder using a separate DT and embed all images in the dataset using the learned vision encoder. Therefore, the data-loading is not bottlenecked by loading entire images into memory, but only their compact representations.
@@
Furthermore, we use global positional embeddings. We also experimented with the Transformer++ recipe (RoPE, SwiGLU, RMSNorm), but only observed minimal performance gains for our problem setting. To speed up training, we use mixed-precision Micikevicius et al. (2017), model compilation as supported in PyTorch (Paszke et al., 2019), and FlashAttention (Dao, 2023).
@@
Implementation. Our implementation of the DT is based on the transformers library (Wolf et al., 2020) and stable-baselines3 (Raffin et al., 2021). We integrated AD, DPT, and RA-DT on top of this implementation.
@@ C.2 Decision Transformer
For Dark-Room and Dark Key-Door, we sample the target return for RTG conditioning before every episode $\mathcal{N}(90,5)$, $\mathcal{N}(370,10)$, and $\mathcal{N}(500,10)$ for grid sizes $10\times 10$, $20\times 20$, and $40\times 20$, respectively. On grid-worlds, we found that sampling the target return performs better than using a fixed target return per grid size. We assume this is because specifying a particular target return biases the DT towards particular goal locations. For MazeRunner, we use a constant target return of 3. For Meta-World, DMControl, and Procgen, we set the target return to the maximum return achieved for a particular task in the training datasets. However, we also found that constant target returns per domain work decently.
@@ C.3 Algorithm Distillation
AD obtains a context trajectory and learns to predict actions of an input trajectory taken $K$ episodes later. Therefore, we tune $K$ per domain. On grid-worlds, we found $K=100$ to perform the best, similar to Lee et al. (2023). For MazeRunner and Meta-World, we set $K=1000$, and for DMControl and Procgen, we set $K=250$.
@@ C.4 Retrieval-Augmented Decision Transformer
Embedding Model. For the embedding model $g(\cdot)$, we either use a DT pre-trained on the same environment with the same hyperparameters as listed in Section C, or a pre-trained and frozen LM. For the pre-trained LM, we use bert-base-uncased from the transformers library by default. BERT is an encoder-only LM with 110M parameters, vocabulary size $v=30522$, and embedding dimension of $d_{\text{LM}}=768$ (Devlin et al., 2019). We apply FrozenHopfield with $\beta=10$ to state, action, reward, and RTG tokens (see Equation 2). To achieve this, we one-hot encode all discrete input tokens, such as actions in Dark-Room/MazeRunner/Procgen or states in Dark-Room, and rewards/RTGs in the sequence before applying the FH. For other tokens, such as continuous states/actions as in Meta-World/DMControl, we directly apply the FH. We evaluate other alternatives for the LM in Appendix E.
@@
0: DT $\pi_{\theta}$, embed model $g$, dataset $\mathcal{D}$, gradient steps NN, context len $C$, batch size $B$, eval frequency $E$, loss function $\mathcal{L}$ (cross-entropy or MSE), evaluate, batch-wise procedures retrieve, reweight, and update
[…]
@@
3:  $\mathcal{I}\leftarrow\mathcal{I}\cup\{(g(\tau_{t-C:t}),\tau_{t-C:t+C})\mid t\in\text{range}(0,|\tau|,C)\}$ $\triangleright$ Add k-v pairs of sub-trjs to $\mathcal{I}$
@@
6:  $\mathbf{b}\sim\mathcal{D}$ where $\mathbf{b}=\{\tau_{j}\ \mid 1\leq j\leq B\}$ $\triangleright$ Sample batch of sub-trjs each of length $C$
7:  $\mathbf{q}=g(\mathbf{b})$ $\triangleright$ Construct queries for all sub-trjs
8:  $\mathcal{R}\leftarrow\text{{retrieve}}(\mathbf{q},\mathcal{I})$ $\triangleright$ Retrieve top-$l$ sub-trjs, Eq. 3
9:  $\mathcal{S}\leftarrow\text{{reweight}}(\mathcal{R})$ $\triangleright$ Re-weight top-$k$ sub-trjs, Eq. 4, 5
10:  $\mathbf{a}=\pi_{\theta}(\cdot\mid\mathbf{b},\{\tau_{\text{ret}}\in\mathcal{S}\})$ $\triangleright$ Predict actions for batch
11:  $\pi_{\theta}\leftarrow\texttt{update}(\pi_{\theta},\mathcal{L},\mathbf{a},\mathbf{b})$ $\triangleright$ Perform gradient step, see Appendix C.1 for $\mathcal{L}$
@@
Constructing queries/keys/values. Regardless of whether $g$ is domain-specific or domain-agnostic, we obtain $C$ embedded tokens after applying $g$ to the input trajectory $\tau_{in}$. Subsequently, we apply mean aggregation over the context length $C$ to obtain the $d_{r}$-dimensional query representation. We experimented with aggregating over all tokens or only tokens of a particular modality (state/action/reward/RTG), and found aggregation over states-only to be most effective (see Appendix E.4). As described in Section 3.2.1, we construct the key-value pairs in our retrieval index by embedding all sub-trajectories in the dataset $\mathcal{D}$ using our embedding model $g$, $\mathcal{K}\times\mathcal{V}=\{(g(\tau_{i,t-C:t}),\tau_{i,t-C:t+C})\mid 1\leq i\leq|\mathcal{D}|\}$. To avoid redundancy, in practice, we construct $H/C$ key-value pairs for a given trajectory $\tau$ with episode length $H$ and sub-sequence length $C$, instead of constructing the key and values for every step $t\in[1,H]$. Note that the values, we store $\tau_{i,t-C:t+C}$, contain both the sub-trajectory itself ($\tau_{i,t-C:t}$) and its continuation ($\tau_{i,t:t+C}$). Similar to Borgeaud et al. (2022), we found this choice important for high performance in RA-DT, because it allows the model to observe how the trajectory may evolve if it predicts a certain action (given that the retrieved context is similar enough).
Vector Index. We use Faiss (Johnson et al., 2019; Douze et al., 2024) to instantiate our vector index $\mathcal{I}$. This allows us to search our vector index in $\mathcal{O}(\log M)$ time using Hierarchical Navigable Small World (HNSW) graphs. Consequently, retrieval-augmentation is highly scalable and commonly employed in LMs with large-scale datasets that contain trillions of tokens (Borgeaud et al., 2022). For the datasets we consider, we found it faster to use a Flat index on the GPU as provided by Faiss instead of using HNSW, because our retrieval datasets are small enough. We use retrieval both during training and during inference. It is, however, possible to pre-compute the retrieved trajectories for $\mathcal{D}$ before the training phase to limit the computational demand of retrieval, as suggested by Borgeaud et al. (2022). During evaluation, we can retrieve after every environment step or only after every $t$ environment steps. Here, $t$ represents a trade-off between inference time and final performance. We use $t=1$ for Dark-Room and Dark Key-Door, and $t=25$ for all other environments (see Appendix E.6 for an ablation on this design choice). For all environments, except for Meta-World and DMControl, we provide a single retrieved sub-trajectory in the agent’s context. For Meta-World and DMControl, we found that providing more than one retrieved sub-trajectory benefits the agent’s performance. Therefore, for these two environments, we retrieve the top-4 sub-trajectories, order them by return achieved in that trajectory, and provide their concatenation as retrieved context for RA-DT.
Reweighting. To implement the reweighting mechanism, as described in Section 3.2.3, we first retrieve the top $l\gg k$ experiences and select the top-$k$ experiences according to their reweighted scores. We set $l=50$ in all our experiments.
@@
Embedding Retrieved Context. After the most similar trajectories have been retrieved, we embed the state/action/reward/RTG tokens with a separate embedding layer (as is done for the regular input sequence) before incorporating them via the CA layers. We also experimented with sharing/detaching the regular embedding layers, but found it most effective to maintain separate ones. Furthermore, we experimented with an additional Transformer-based encoder for the retrieved sequences, as proposed by Borgeaud et al. (2022), but did not observe substantial performance gains despite increased computational cost.
One simple strategy to mitigate this issue is deduplication, i.e., to discard duplicate experiences before the training phase of RA-DT. To achieve this, we first construct our index as described in Section 3.2. For every key $\mathbf{k}\in\mathbf{K}$, we retrieve the top-$k$ neighbours (excluding experiences from the same episode as $\mathbf{k}$). If the similarity score is above a cosine similarity of $0.98$, we discard the experience. This substantially reduces the number of experiences in the index and speeds up retrieval. Note that deduplication can be used at inference time to avoid storing redundant experiences.
(a) Dark-Room $\times10$
(b) Dark-Room $\times20$
[…]
Two other strategies for regularizing retrieval during the training phase are similarity cut-off and query dropout (Yasunaga et al., 2023). Similarity cut-off first retrieves the top $m>l$ experiences, discards the experiences with a similarity score above a threshold (e.g., $0.98$), and retains only the remaining experiences $l$. If used in combination with reweighting, we set m=2lm=2*l. Query dropout randomly drops out tokens (e.g., 20%) of the embedded sub-trajectory $\tau_{\text{in}}$, which leads to more diverse retrieved experiences. We found both strategies effective for RA-DT. We use a query dropout of 0.2, a similarity cut-off of 0.98, and deduplication by default. Furthermore, for Meta-World and DMControl, we found query-blending useful. Query-blending interpolates between the actual query and a randomly selected key from the retrieval index, $\bm{q}^{\prime}=\bm{q}*\alpha_{\text{blend}}+(1-\alpha)\bm{q}_{\text{rand}}$. For Meta-World and DMControl, we additionally set $\alpha_{\text{blend}}=0.5$.
On Dark-Room and Dark Key-Door environments, we found it useful to replace retrieved experiences with experiences randomly sampled from the same task if the query sub-sequence is from the beginning of the episode (i.e., smaller than timestep 10). This is because in these two environments, retrieving appropriate experience can be difficult if the given query sub-sequence is too short.
(a) 80 Train Goals
(b) 20 Eval Goals
Figure 15: Average performances on Dark-Room $\times10$ over the course of training for (a) train and (b) test tasks. We train each agent for 100K steps and evaluate every 25K steps. Curves are averaged across the 80 training and 20 evaluation tasks, respectively. We report mean reward (+ 95% CI) over 3 seeds.
Figure 16: Attention map analysis for an optimal context-trajectory on Dark-Room $10\times 10$. We plot the retrieved context trajectory (left), the corresponding attention map, and the actual agent state (right), across timesteps (1, 5, 10). Queries (input trajectory) are on the y-axis and keys (context trajectory) on the x-axis. We highlight the sub-sequence in the context trajectory with the highest attention score (left). To improve readability, we mask out attention scores below a certain threshold and only provide labels for tokens that exhibit the highest attention scores. The agent imitates the context trajectory and successfully finds the goal.
Figure 17: Attention map analysis for a suboptimal context-trajectory on Dark-Room $10\times 10$. The agent selects a different route than is present in the suboptimal context trajectory and explores the environment.